1.CALCULATE THE NO OF WAVES MADE BY ELECTRON IN 5TH BOHR ORBIT OF H ATOM
2.A HYDROGEN LIKE SYSTEM HAS IONIZATION ENERGY 11808 KJ/MOL.FIND NO OF PROTONS IN NUCLEUS OF ATOM

Grade:Upto college level

2 Answers

Chetan Mandayam Nayakar
12 years ago

mvr=5h/2pi,pr=5h/2pi, p=5h/2pir,lambda= h/p=2pir/5 thus number of waves made by electron=5, let z be number of protons,

ionization energy =13.6*1.6*(10^-19)*6.02*(10^23)*z2=11808000, solve for z

 

SAGAR SINGH - IIT DELHI
12 years ago

Dear student,

No. of waves=n(n-1)/2

BEST OF LUCK..!!!!


 

We are all IITians and here to help you in your IIT JEE preparation.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar : Click here to download the toolbar..

 

Askiitians Expert

Sagar Singh - IIT Delhi